How many molecules or SO2 are produced from 13.1 grams of CS2

1
Expert's answer
2021-01-20T03:17:29-0500

CS2+3O2 = CO2 + 2SO2

13.1 x 6.02 x 10^23 x2/76= 2,07531579e23 .